‏Its hunting season and Natalia wants part of the action so she grabs her Nerf gun and fires a 2 gram foam dart [ W 42 ° N ] into the air with a velocity of 24m / s . A goose , travelling at 5m / s [ W ] gets pegged and it isn't impressed , so it drops a payload of goose poop on Natalia's shiny Mercedes convertible . ( Yup ... the top was down ) What height did the Nerf dart reach ? ( Assume Natalia stands 1.6m tall ) How long does it take for the goose's payload to hit the Mercedes ? ( Assume the Mercedes is 1.6m tall )


1
Expert's answer
2021-10-12T12:30:40-0400

"m= 2g = 0.002\textsf{ g}\\
v= 24\textsf{ m/s}\\
u_y = 24\sin 42° = 16.06\textsf{ m/s}\\
v_y² = u_y² -2gh\\
0² = 16.06²-2(9.8)h\\
h = 16.06²/2(9.8) = 13.15m"


Total height = Height of projectile + Natalia's height = 13.15 + 1.6 = 14.75m"."